OpenStudy (gabylovesyou):

Create your own unique quadratic equation in the form y = ax2 + bx + c. Use complete sentences and show all work to determine the following: Does the graph open up or down? How do you know? (2 points) Explain whether the graph has a maximum or minimum point. (2 points) Find the vertex and x-intercepts of the graph. (3 points)

OpenStudy (anonymous):

Can you start ?

OpenStudy (gabylovesyou):

would this be an example? x^2 + 6x + 4

OpenStudy (anonymous):

Sure ,why not !! Now Can you determine the following : Does the graph open up or down? How do you know? (2 points) Explain whether the graph has a maximum or minimum point. (2 points) Find the vertex and x-intercepts of the graph. (3 points)

OpenStudy (gabylovesyou):

the graph opens up right?

OpenStudy (anonymous):

True ,Since (a) is +ve :)

OpenStudy (gabylovesyou):

the graph has a minimum point?

OpenStudy (anonymous):

Since it opens upwards, it must have a minimum value. Graph it and find it out :)

OpenStudy (gabylovesyou):

yea its minimum :) what is the formula for vertex?

OpenStudy (gabylovesyou):

@Eyad

OpenStudy (anonymous):

Since this is a quadratic equation , Vertex = b/2a

OpenStudy (phi):

you chose would this be an example? x^2 + 6x + 4 this is a valid answer, but it is "hard" to find its x-intercepts (you have to use the quadratic formula because it does not factor nicely. If you start with for example, (x+1)(x-2) and multiply it out you get a nice quadratic

OpenStudy (gabylovesyou):

:o i want the easiest way lol so let me do (x + 1)(x - 2)

OpenStudy (gabylovesyou):

how do i solve that ^^^^^^^^^^^^^^^ ???

OpenStudy (phi):

use FOIL to multiply it out so it is in "standard form" can you mutliply it out ?

OpenStudy (gabylovesyou):

x^2 .... ummmm

OpenStudy (phi):

FOIL is short for First x*x outer -2*x inner 1*x last -2*1 you get x^2 -2x + x -2 combine the -2x+1x to get -1x finally: y= x^2 -x -2 remember it factors into (x+1)(x-2) and you can solve for (x+1)(x-2)=0 to find the x-intercepts

OpenStudy (phi):

this makes a smile because the "a" is positive (a is the number in front of x^2, a is +1)

OpenStudy (gabylovesyou):

so it still opens up and it still has a minimum point

OpenStudy (gabylovesyou):

@phi how do i find the vertex?

OpenStudy (gabylovesyou):

@phi ;/

OpenStudy (phi):

your quadratic is y= x^2 -x -2 this is in the form a x^2 + bx +c where a =1, b= -1 and c= -2 the vertex occurs at x = -b/(2a) replace the letters with the numbers. what do you get for x ?

OpenStudy (gabylovesyou):

-1/2*1 ????

OpenStudy (gabylovesyou):

- 1/2 ??

OpenStudy (phi):

what is -b ? it is - (-1) and the 2 minus signs make a + try again

OpenStudy (gabylovesyou):

so it would be - (-1)/2*1 ?

OpenStudy (phi):

yes though you should write it as - (-1)/(2*1) with the 2*1 in parens... that simplifies to what ?

OpenStudy (gabylovesyou):

1/2

OpenStudy (gabylovesyou):

right?

OpenStudy (phi):

yes, x= 1/2 now find the y value that goes with this x value replace x with 1/2 in y =x^2 -x -2 you can use a calculator if you have one

OpenStudy (gabylovesyou):

wait.. the vertex was x = 1/2 ? or we r in the process ?

OpenStudy (phi):

yes the vertex has an x value of 1/2. I assume they want the (x,y) pair for the vertex, so you should find the y value also

OpenStudy (gabylovesyou):

ok.. y = -9/4

OpenStudy (phi):

yes, so the vertex is at (0.5, -2.25) in decimals or (1/2, -9/4) as fractions finally, the x-intercepts happen when (x+1)(x-2) =0 this means either x+1 =0 or x-2=0 solve both equations to find the 2 x-intercepts

OpenStudy (gabylovesyou):

do u normally do it fractions or decimals?

OpenStudy (phi):

as a check, we can use wolfram http://www.wolframalpha.com/input/?i=plot+y%3D%28x%2B1%29%28x-2%29+ click on the properties button in the bottom panel to see the vertex fractions are fine. In the real world, we use decimals, but in school fractions are easy to read

OpenStudy (phi):

to solve x+1=0 you could add -1 to both sides: x + 1 -1 = 0 - 1 and simplify to get x = -1 I assume you can solve x-2 = 0 ?

OpenStudy (gabylovesyou):

now we r gonna find the x- intercepts right?

OpenStudy (gabylovesyou):

@phi

OpenStudy (phi):

yes, find the x-intercepts. but because we started with y= (x+1)(x-2) you can find them by solving (x+1)(x-2) = 0 (x intercepts are the x values that make y = 0) to solve x+1=0 you could add -1 to both sides: x + 1 -1 = 0 - 1 and simplify to get x = -1 I assume you can solve x-2 = 0 ?

OpenStudy (gabylovesyou):

x = 2

OpenStudy (gabylovesyou):

@phi

OpenStudy (phi):

I think you have answered all the questions. the x-intercepts can also be written as (x,y) pairs: (-1,0) and (2,0) but it is generally understood that the y value of an x-intercept is 0

OpenStudy (gabylovesyou):

so the x intercepts are (-1, 2)

OpenStudy (gabylovesyou):

thank you!!

OpenStudy (phi):

I would not put them in parens, because that might be confused with the point (-1,2) best to say (-1,0) and (+2,0)

OpenStudy (gabylovesyou):

ok :)

OpenStudy (anonymous):

let y=2x^2+4x+3 =2(x^2+2x)+3 =2(x^2+2x+1-1)+3 =2(x+1)^2 -2+3 =2(x+1)^2+1 or (y-1)=2(x+1)^2 or (x+1)^2=1/2 (y-1) It is an upper parabola with vertex at (-1,1) &x-intercept of vertex=-1